OpenStudy (anonymous):

Solve for x. 8x - 4 = 2x - 3

OpenStudy (anonymous):

Follow similar methodolgy like in the previous problem. Give it a try on your own first :)

OpenStudy (anonymous):

\[8x-2x=-3+4 ...................so.......6x=1.......and.......x=1/6\]

OpenStudy (anonymous):

8x - 4 = 2x - 3 8x - 2x = -3 + 4 6x = 1 x = 1/6

OpenStudy (anonymous):

x=1/6